OpenStudy (anonymous):

At the airport, you pull a 13kg suitcase across the floor with a strap that is at an angle of 49∘ above the horizontal. PLEASE SOMEONE HELP!

OpenStudy (shamim):

Check ur question once again!!!!!

OpenStudy (anonymous):

what do you mean?

OpenStudy (shamim):

Wt do u wanna know?

OpenStudy (shamim):

But i can say something abt ur question

OpenStudy (shamim):

U know the weight of 13kg is 13*9.8N acting downward

OpenStudy (shamim):

If u apply F force Then the horizontal component of F is=F*cos49

OpenStudy (shamim):

Vertical component of F=F*sin49

OpenStudy (shamim):

13*9.8N and F*sin49 are oposit in direction but not equal

OpenStudy (shamim):

So net force acting downward =13*9.8-F*sin49

OpenStudy (shamim):

Is it clear wt i m telling???

OpenStudy (anonymous):

Yes, thank you.